OpenStudy (anonymous):

Nicole missed the lesson on normal distribution and needs to do her homework. Explain to Nicole how to use the mean and standard deviation of a normal distribution to determine the top 5% of the population.

OpenStudy (sidsiddhartha):

ok do u have the z-chart with u?

OpenStudy (anonymous):

whoops, z chart?

OpenStudy (sidsiddhartha):

here \[\phi[\frac{ x-\mu }{ \sigma }]=1-0.05=0.95\] so \[\frac{ x-\mu }{ \sigma }=\phi(0.95)\] so u need to substitute the z-score of 0.95 here

OpenStudy (anonymous):

oh okay so that's all i have to do?

OpenStudy (sidsiddhartha):

yeah that'll be all after that u just have to put the values of \[\mu ,\sigma\]

OpenStudy (anonymous):

so what formula is that called?

OpenStudy (sidsiddhartha):

uhh u can call it normal distibution formula where\[\mu=mean\]and \[\sigma=standard. deviation\]

OpenStudy (anonymous):

oh thanks so much

OpenStudy (sidsiddhartha):

np :)